OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9

OpenStudy (anonymous):

not clear ya

OpenStudy (anonymous):

3(y + 2) + 2y = 9 3(–y + 2) + 2y = 9 3x + 2(y + 2) = 9 3x + 2(–y + 2) = 9